OpenStudy (anonymous):

Explain how the formula for the area of a square can be derived from the formula for the area of a rectangle?

OpenStudy (epoweritheta):

good question ; A=lb for any rectangle now , consider the rectangle which has l=b ; (which is also a square) A=l*l=l^2

OpenStudy (anonymous):

Thank You that helped